OpenStudy (anonymous):

3(t + 2) + t = 2(3 + 2t)

OpenStudy (anonymous):

what the question wants ? find t or prove both sides are the same ?

OpenStudy (anonymous):

Find T

OpenStudy (phantom):

t should be then any real number as the equation is an identity, i.e true for all values of t

OpenStudy (phantom):

If you are trying to solve in the complex domain, then t can be any complex number too

OpenStudy (anonymous):

t=3 3Xt+2= 3X3+2=15 15+t= 15+3=18 = 2X3+2t= 2X3+(2X3)= 2X3+6= 2X9= 18 Before brackets means you times that number by whatever results you get within the bracket. So for the first side of the equation, it's 3 times the bracket. The bracket is t+2 which is 3+2 which results in 5 which times by 3 (the number before the bracket) is 15 and then you take what's after the bracket (in this case +t) which results in 15+3 which is 18. And the same process for the other side which ends you up with 18 as well. c/p from yahoo answer

OpenStudy (phantom):

check for t=4 ---------- LHS : 3(t+2) +t = 22 RHS: 2(3+2t) = 22 again LOL..thats why I said its an identity !!

OpenStudy (anonymous):

yess, its like trying an error. any number that will satisfy both sides of this equations.

Directrix (directrix):

If you are trying to solve in the complex domain... ------------------- All Complex Numbers